Shamrock Co. sells $365,000 of 12% bonds on June 1, 2020. The bonds pay interest on December 1 and June 1. The due date of the bonds is June 1, 2024. The bonds yield 8%. On October 1, 2021, Shamrock buys back $116,800 worth of bonds for $123,800 (includes accrued interest). Give entries through December 1, 2022.(b) Shamrock Co. sells $365,000 of 12% bonds on June 1, 2020. The bonds pay interest on December 1 and June 1. The due date oPrepare all of the relevant journal entries from the time of sale until December 31, 2022. (Assume that no reversing entries(To record buy back of bonds) 12/1/21 12/31/21 6/1/22 12/1/22
Answer #1
(b) Schedule of bond premium amortization
Effective interest method
Bonds sold to yield
Date Cash paid Interest expense Premium amortized Carrying amount of bonds
a b c=a-b
6/1/20. 414149
12/1/20. 21900 16566 5334 408815
6/1/21. 21900 16353 5547 403268
12/1/21. 21900 16131 5769 397498
6/1/22. 21900 15900 6000 391498
12/1/22. 21900 15660 6240 385258
6/1/23. 21900 15410 6490 378768
12/1/23. 21900 15151 6749 372019
6/1/24. 21900 14881 7019 365000
Interest expense=Beginning carrying value of the bond*Bonds yield for 6 months
Carrying value of the bond=Beginning carrying value of the bond-Premium amortized
Date Account titles and explanation Debit Credit
June 1,2020. Cash (Note:1) 414149
Premium on bonds payable (Plug) 49149
Bonds payable 365000
(Issue of bonds at premium)
Dec 1,2020. Interest expense (Ref amortization schedule) 16566
Premium on bonds payable 5334
Cash 21900
(Interest paid)
Dec 31,2020. Interest expense 16353*(1/6) 2725
Premium on bonds payable (Plug) 925
Interest payable 21900*(1/6) 3650
(Interest accrued for Dec month)
June 1,2021. Interest expense 16353*(5/6) 13628
Interest payable 3650
Premium on bonds payable (Plug) 4622
Cash 21900
(Interest paid)
Oct 1,2021. Interest expense (Note:2) 3441
Premium on bonds payable (Plug) 1231
Cash 4672
(Interest paid on bonds bought back)
Oct 1,2021. Bonds payable 116800
Premium on bonds payable (Note:3) 11852
Gain on redemption of bonds (Plug) 9524
Cash (By reducing interest) (123800-4672) 119128
(To record buy back of bonds)
Dec 1,2021. Interest expense (16131*68%) 10969
Premium on bonds payable (Plug) 3923
Cash (21900*68%) 14892
(Interest paid)
Dec 31,2021. Interest expense 15900*68%*(1/6) 1802
Premium on bonds payable (Plug) 680
Interest payable 21900*68%*(1/6) 2482
(Interest accrued for Dec month)

9010 2482 3400 14892 June 1,2022. Interest expense 15900*68% *(5/6) Interest payable Premium on bonds payable (Plug) Cash (21Note:2 % of bond bought back=116800/365000=0.32=32% Interest expense for 4 months=16131*32%*4/6=$ 3441 Interest paid for 4 mo